OpenStudy (anonymous):

What kind of angle is formed when two right angles are placed back to back, so they have one ray in common? The options are: A.) Acute B.) Obtuse C.) Right D.) Straight

OpenStudy (anonymous):



Directrix (directrix):

Check out the attached chart. Post what you think the answer is. Somebody will be by to check. Okay? http://www.mathsisfun.com/angles.html

OpenStudy (anonymous):

Okay, I think that the answer would be C.) RIght...

Directrix (directrix):

You have two right angles that are adjacent. So 90 + 90 = ?

OpenStudy (anonymous):

Ohh so it would be Obtuse, because 90+90=180, and and Obtuse angle is Between 90 and 180 degrees...?

Directrix (directrix):

No. 90 + 90 = 180 which is not obtuse. Look at the chart.

Directrix (directrix):